A particle starts from rest and moves in a straight line on a smooth horizontal surface. Its acceleration at time t seconds is k(4v + 1)m/s^2 , where k is a positive constant and v m/s is the speed of the particle. Given that v = (e^2 - 1)/4 when t = 1, show that v = 1/4(e^(2t) - 1)
1
Expert's answer
2021-02-07T19:18:34-0500

A particle starts from rest  and the initial velocity is 0

"a = \large\frac{v - v_o}{t} \to v = at"

"a = k(4v+1) \large\frac{m}{s^2}"

Given that "v = \large\frac{e^2-1}{4}" when t = 1

"v(1) = a =" "\large\frac{e^2-1}{4}" "= k(4v+1)"

"\large\frac{e-1}{2}*\frac{e+1}{2}" "= k * (4v+1)"

"v = \large\frac{e^{2t} - 1}{4}" "\to v(1) = \large\frac{e^2-1}{4}"